OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {14, 18, 21, 15, 17} Data Set 2: {15, 17, 22, 20, 16}

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

its it 0.2

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

*is it

jigglypuff314 (jigglypuff314):

I don't believe 0.2 is a ratio? what did you get as the mean for the first set? :)

OpenStudy (anonymous):

17

OpenStudy (anonymous):

@jigglypuff314 i got 17 for the first set

jigglypuff314 (jigglypuff314):

that's right :) do you know how to find the MAD?

OpenStudy (anonymous):

umm nope

jigglypuff314 (jigglypuff314):

you basically find the difference between each value and the Mean so for Set 1 you would get 17 - 14 = 3 18-17 = 1 etc {3, 1, 4, 2, 0} would be the MAD set, then find the Mean of that :)

OpenStudy (anonymous):

okay can stay so you can see if i get something right?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (anonymous):

the mean of the second one is 18

jigglypuff314 (jigglypuff314):

yes ^_^

OpenStudy (anonymous):

so then i subtract

jigglypuff314 (jigglypuff314):

mhmm

OpenStudy (anonymous):

15-18=3

jigglypuff314 (jigglypuff314):

you're doing great :) keep going

OpenStudy (anonymous):

17-18=1

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

3,1,i need some help

OpenStudy (anonymous):

{3, 1, 4, 2, 0} is it the same as this ?

jigglypuff314 (jigglypuff314):

not quite you need to keep going 22 - 18 = 4 20 - 18 = 2 16 - 18 = 2 so {3, 1, 4, 2, 2}

OpenStudy (anonymous):

so find the mean then that will be my answer?

OpenStudy (anonymous):

the mean it 2.4

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

that's right :)

OpenStudy (anonymous):

so 2.4 is the answer ?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

not the answer it's part of it ratio mean : MAD

OpenStudy (anonymous):

oh i need to round right?

jigglypuff314 (jigglypuff314):

no, just put 18 as the mean and 2.4 as the MAD

jigglypuff314 (jigglypuff314):

18:2.4